OpenStudy (anonymous):

Pamela is 8 years younger than Jiri. The sum of their ages is 84 . What is Jiri's age?

OpenStudy (anonymous):

50

OpenStudy (anonymous):

never mind sorry wrong question

OpenStudy (anonymous):

the answer to Jiri's age is 46

OpenStudy (anonymous):

P=J-8 P + J = 84 J-8+J =84 2J-8=84 2J=92 J=46